# Low Fidelity Visuo-Tactile Pretraining Improves Vision-Only Manipulation Performance

### v3

@@ Abstract
Tactile perception is a critical component of solving real-world manipulation tasks, but tactile sensors for manipulation have barriers to use such as fragility and cost. In this work, we engage a robust, low-cost tactile sensor, BeadSight, as an alternative to precise pre-calibrated sensors for a pretraining approach to manipulation. We show that tactile pretraining, even with a low-fidelity sensor as BeadSight, can improve an imitation learning agent’s performance on complex manipulation tasks. We demonstrate this method against a baseline USB cable plugging task, previously achieved with a much higher precision GelSight sensor as the tactile input to pretraining. Our best BeadSight pretrained visuo-tactile agent completed the task with 70% accuracy compared to 85% for the best GelSight pretrained visuo-tactile agent, with vision-only inference for both. Code for this project is available at: https://github.com/selamie/beadsight
Keywords: Tactile sensing, robotic manipulation, contrastive pretraining, imitation learning
Figure 1: Our method uses a low-fidelity tactile sensor for pretraining. Contrastive tactile-vision observation pretraining produces a visual encoder well-aligned to tactile observations that is then fine-tuned on the actual action sequence prediction task. Finally, vision-only performance at inference is improved using the residual tactile influence of the pretrained encoder.
1 Introduction
@@
Translating advances in visual perception to robotic grasping and manipulation of objects remains challenging. For complex manipulation tasks such as peg insertion, pulling or twisting with resistance, and dynamic motions such as throwing and catching, fine-grained manipulation requires tactile perception. Tactile sensors have been paired with visual sensors for both classical control and machine learning approaches to these tasks [1], but issues of fragility and cost present barriers to heavy use or industrial integration, particularly for manipulation tasks that would place higher forces on sensors at the tactile edge.
@@
Previously, a GelSight [2] tactile sensor was used to train an agent on a USB insertion task [3], the first time this was achieved with imitation learning. GelSight is not designed for robustness to higher shear forces and was noted to break irrecoverably during data collection and inference for that task, requiring repeated replacement. This work also demonstrated an approach using tactile information only during pretraining, then ablating the tactile sensor at inference, achieving a more robust vision-only manipulation system.
@@
BeadSight [4] aimed to make a simpler, low cost calibration-free sensor that, like GelSight, still operated at an end effector’s point of contact with objects. We constructed the BeadSight sensor, which does not rely on any calibration and instead relies entirely on neural networks to distill information about contacts and movements at the tactile edge.
In this work, we repeated the visuo-tactile pretraining USB plugging experiment using the much lower fidelity BeadSight to produce a direct comparison with the GelSight sensor in the task of plugging in a USB cable. Our contributions are:
@@
    We demonstrate the effectiveness of visuo-tactile pretraining even with a lower fidelity, imprecise sensor on two different state-of-the-art imitation learning methods.
@@
    We demonstrate an affordable durable approach to visuo-tactile pretraining, which is especially cost-effective for improving the performance of vision-only agents.
@@
    We show that by freezing a pretrained tactile encoder, we can mitigate the overfitting of low-fidelity tactile sensors.
2 Background and Related Work
2.1 Tactile Sensing
In robotic manipulation, there are a wide variety of tactile sensors and sensing approaches [1]. These sensors use different modalities to collect tactile information, including forces and torques applied normally or tangentially [5] [6], mechanical vibrations [7] [8], thermal measurement and conductivity [9] [10], and pre-touch proximity [11].
@@ 2.1 Tactile Sensing
BeadSight is a low-cost visuo-tactile sensor that uses hydrogel beads as an inexpensive, durable sensing medium [3]. This deformable surface is inexpensive and durable, with an embedded camera observing the beads’ motion and deformation. However, the beads’ movements are stochastic and, unlike GelSight, the relationship between the beads and shear forces cannot be directly resolved with analytical methods. Previously, a deep neural network was used to reconstruct contact pressure maps with BeadSight as an example tactile learning approach.
2.2 Tactile Sensing and Manipulation Control Policies
Prior work integrating tactile sensors in control policies range from classical control approaches to reinforcement and imitation learning. [16] used classical state estimation and a heuristic method to follow cables and insert wires. [17] combined a linear-quadratic regulator (LQR) controller and linear dynamics model to follow an audio cable and successfully plug in a headphone jack. Reinforcement Learning (RL) agents have been trained to perform peg insertion, door opening, and in-hand rotation tasks [18, 19]. These approaches require simulation for the sheer amount of training required ($\sim$ 1 million steps). Nearest Neighbors Imitation Learning [20] is a method that encodes observations and demonstrations into a latent space and compares the distance between them; [21] used this method with online residual reinforcement learning to reduce computation intensity and learn visuo-tactile peg insertion. [22] used a similar approach to learn a variety of tasks including bowl and cup unstacking, bottle opening, and joystick movement.
2.3 Imitation Learning Methods
Imitation Learning (IL) methods aim to mimic human behavior by learning from human demonstrations, which can be easier than designing optimal reward functions as in reinforcement learning [23]. Variations of this method have trained manipulation agents to perform kitchen tasks and stack blocks [24, 25] and even sculpt clay [26]. IL tends to perform more poorly for more complex action sequences, particularly with non-deterministic goal policies and significant shifts between training and deployment domains. We focus on state of the art methods that aim to address these drawbacks, Action Chunking Transformers (ACT) [27] and Diffusion Policy [28]. ACT uses a temporal ensembling approach to maintain a consistent action trajectory in spite of out-of-distribution observations, and Diffusion Policy uses diffusion to better model multi-modal action pathways.
2.3.1 Diffusion Policy
@@
2.3.2 Action Chunking Transformers (ACT)
@@
ACT trains a Conditional Variational Autoencoder (CVAE) combined with a transformer backbone to predict action sequences conditioned on state and vision observations [27]. During training, the latent variable of the CVAE helps the model capture multi-modal policies, while a large KL divergence loss keeps the network from becoming too reliant on the latent information. During inference, the latent viable is set to zero. The autoencoder uses a latent variable to reduce the impact of multi-modal training data, and temporal ensembling at inference reduces the effect of a single bad prediction. Specifically, ACT predicts the goal states $A_{t}$ to $A_{t+h}$ at each timestep. During inference, the previous predictions $A_{t-h}$ to $A_{t-1}$ are combined with current predictions using a weighted average of exponentially decaying weights, $w_{i}=e^{-ki}$. These “ensembled” actions are then executed.
2.4 Contrastive Pretraining
@@
Contrastive pretraining methods train models to distinguish between similar and dissimilar data, coding them as positive (similar) or negative (dissimilar) pairs [31]. Those modeled after the structure of Contrastive Language Image Pretraining (CLIP) [32] do this by maximizing the cosine similarity of the paired data in some multimodal embedding space, while minimizing the similarity of dissimilar pairs. For visuo-tactile robotic manipulation, tactile and visual observations can be compared in that latent space. [33] used this approach to better identify flaws in fabric, and [34, 35] both improved visuo-tactile object identification this way.
3 Methods
Figure 2: Left, BeadSight components and construction. Slight modifications were made to the original enclosure to better manipulate objects, including the robot gripper mount and a separate housing lid to create a secure, interchangeable deformable surface. Right, the embedded camera’s view of the BeadSight hydro-gel bead sac, relaxed (1) and being pressed (2), after down-scaling and correcting fish eye distortion.
3.1 BeadSight Hardware
@@
The version of BeadSight used for this work was constructed in a multi-part design that separated the deformed surface, camera, and contact surface. The BeadSight enclosure is a 3D printed body with threaded heat inserts on attachment points, and includes a mount that attaches the body to the Franka Emika Panda robot gripper. A ”lid” allows easy separation of an interchangeable contact surface, making the system repairable for long-term wear and tear. The lid consists of a 3D printed frame, press-fit acrylic insert, and a hydro-gel bead sac. The hydro-gel bead sac is backed with white adhesive tape to reflect light back to the camera and improve visibility of the deformation. Black Polyvinyl Chloride (PVC) adhesive tape is the final layer on the hydro-gel bead sac, sealing the sensor from external light sources.
@@
The hydro-gel bead sacs themselves are of the same construction as [4], using Polyacrylamide (PAM) beads that are hydrated with water and sealed in clear PVC sheeting. For the data in this work, the bead sac used about 50 green PAM beads hydrated with 5ml of water. The embedded camera is also the same as in [4], a 1080 $\times$ 1920 USB camera with 180°fish eye lens. 4 LEDs illuminate the hydro-gel sac, with the camera capturing at 30Hz. The image frames from the camera are down-scaled to 480 $\times$ 480 resolution and the fish-eye is corrected before processing.
3.2 Data Collection
Figure 3: Experimental scene, with a D415 wrist camera (not visible) behind the robot gripper.
Our experimental setup recreated [3] in order to directly compare the performance of BeadSight and GelSight. To train ACT, Diffusion Policy, and the contrastive pretraining model, we used 105 expert demonstrations collected via teleoperation of a Franka Emika Panda robot. We used the Oculus Virtual Reality (VR) teleoperation method developed in [36] which uses handheld Oculus controllers to move the robot end-effector. Only motion control was done with the Oculus system; the human teleoperator looked directly at the workspace rather than through the VR headset.
@@ 3.2 Data Collection
Visual observations of the scene were collected by six Intel RealSense scene cameras. These consist of four D415 cameras mounted around the workspace, a D415 camera mounted at the robot’s wrist, and a D445 overview camera mounted on a top edge of the workspace frame. The BeadSight sensor mounted on the robot gripper fingers collected tactile observations. A parallel finger was 3D printed to match the BeadSight, with silicone gel on the surface to aid grip.
The experimental task consisted of a USB port hub, a holder for a USB plug, and the USB plug with friction tape applied to assist grip. The hub and holder were fixed for the duration of the experiments, as our goal was to evaluate the plug insertion itself rather than the ability to locate the plug or port. We collected 105 expert demonstrations of grasping the plug and inserting it into the USB port closest to the robot (a successful execution at inference must choose the same port).
3.3 Tactile Encoder
@@
The hydro-gel beads are not fixed components, and therefore cannot record any exact tactile measurements. Instead, tactile events can be inferred from the beads’ relative motion and deformation over some period of time. To capture this relative motion, we introduce a “tactile horizon”, the total number of past tactile frames (including the current frame) combined to form the tactile observation. We use a tactile horizon of 5 frames (the current time step and four previous time steps). We collapse the 5 frames on the channel dimension, converting 5 individual $3\times 480\times 480$ images to a single $15\times 480\times 480$ tensor. For a different tactile horizon, the representative tensor would be of the shape $3h\times 480\times 480$, where $h$ is the horizon. This collapsed tensor allows us to add a single convolutional layer (converting 15 channels to 3 channels) to a pre-trained ResNet-18 [37] encoder as the encoder for BeadSight. The classification layer is removed such that a given image is projected to a 512 embedding vector before being passed to the downstream model.
@@
To provide a horizon for the first few time steps, the BeadSight tactile observation buffer is initialized with $h-1$ duplicates of the initial tactile frame. For this task, the sensor is not in contact with any object or surface at the start of each task run, so we would expect the initial frames to be nearly identical.
3.3.1 Contrastive Pretraining
@@
The vision and tactile encoders were pre-trained using the contrastive pretraining method described in [3]. This method leverages the relationships between the three different modalities of data our agent observes (visual, tactile, and positional) to train the encoders to extract task-relevant features. During pretraining, two projection networks are learned, one for visual observations, and one for tactile and positional observations. These projection networks consist of an encoder (ResNet-18 for the image encoder, and the BeadSight encoder described above for the tactile encoder), along with a projection head that casts the encoded observations to a shared latent embedding space. The positional information is also passed to the tactile projection head. By adding position information to the tactile projector, we ensure the resulting latent space has both global information (position observations) and local information (tactile observations).
@@
To train the projection networks, a set of timesteps is sampled from a single trajectory, and the corresponding observations are projected to the shared latent space. Then, a CLIP loss is used to encourage the embeddings from the same timestep to be similar, while driving the projection of the observations from different timesteps. All of the camera views are encoded using a shared visual encoder. At each update, the constrastive loss is calculated separately for each camera and then summed, according to the equation:
$loss=-\sum_{c}\sum_{i\in t}\frac{1}{2n}(log(\frac{exp(sim_{i,i,c}/\tau)}{\sum_{j\in t}exp(sim_{i,j,c}/\tau)})+log(\frac{exp(sim_{i,i,c}/\tau)}{\sum_{j\in t}exp(sim_{j,i,c}/\tau)}))$
@@
3.3.2 Diffusion Policy
Our Diffusion Policy approach leveraged [28], and was modeled after [3], with similar baseline hyper-parameters. We used the 1-D temporal CNN implementation which was noted to be less sensitive to hyper-parameter tuning and train stably for most tasks. Diffusion Policy is most effective when each camera view uses a separate initial encoder, so 6 ResNet-18 encoders were constructed and fine-tuned during training, one for each scene camera. We remove the final ResNet-18 layer to project the image to a 512 embedding for the downstream model. For the pretrained model, six of the pretrained ResNet-18 vision encoders were constructed and separately fine-tuned during training.
After encoding, the vision and tactile observations were stacked to form a single observation vector, then passed to the network to condition its noise prediction. To compare directly to GelSight’s performance in [3], we use the same settings of 1 observation horizon, an action prediction horizon of 20, and executed 8 timesteps out of the 20 available predictions at inference. We also took advantage of noise scheduler decoupling, using 100 denoising steps during training for better resolution but using only 10 inference steps for faster execution.
3.3.3 Action Chunking Transformers

### v4

# Low-Fidelity Visuo-Tactile Pre-Training Improves Vision-Only Manipulation Performance

@@ Abstract
Tactile perception is essential for real-world manipulation tasks, yet the high cost and fragility of tactile sensors can limit their practicality. In this work, we explore BeadSight (a low-cost, open-source tactile sensor) alongside a tactile pre-training approach, an alternative method to precise, pre-calibrated sensors. By pre-training with the tactile sensor and then disabling it during downstream tasks, we aim to enhance robustness and reduce costs in manipulation systems. We investigate whether tactile pre-training, even with a low-fidelity sensor like BeadSight, can improve the performance of an imitation learning agent on complex manipulation tasks. Through visuo-tactile pre-training on both similar and dissimilar tasks, we analyze its impact on a longer-horizon downstream task. Our experiments show that visuo-tactile pre-training improved performance on a USB cable plugging task by up to 65% with vision-only inference. Additionally, on a longer-horizon drawer pick-and-place task, pre-training — whether on a similar, dissimilar, or identical task — consistently improved performance, highlighting the potential for a large-scale visuo-tactile pre-trained encoder. Code for this project is available at: https://github.com/selamie/beadsight.
I INTRODUCTION
Figure 1: A diagram of the flow of information through tactile pre-training and inference. During contrastive pre-training, primed tactile and visual encoders are produced. The encoders are further refined on the downstream task. Then, the tactile encoder is discarded for vision-only inference without the need for a tactile sensor.
@@ I INTRODUCTION
Translating advances in visual perception to robotic grasping and manipulation of objects remains challenging. For complex manipulation tasks such as peg insertion, pulling or twisting with resistance, and dynamic motions such as throwing and catching, fine-grained manipulation requires tactile perception. Tactile sensors have been paired with visual sensors for both classical control and machine learning approaches to these tasks [1], but fragility and cost issues present barriers to heavy use or industrial integration, particularly for manipulation tasks that would place higher forces on these sensors.
@@
Previously, a GelSight [2] tactile sensor was used to train an agent on a USB insertion task [3]. GelSight is not designed for robustness to higher shear forces and was noted to break irrecoverably during data collection and inference for that task, requiring repeated replacement. This work also demonstrated an approach using tactile information only during pre-training, then disabling the tactile sensor for fine-tuning and inference, achieving a more robust vision-only manipulation system.
@@
BeadSight [4] aimed to make a simpler low-cost, high-durability sensor that, like GelSight, still operated at an end effector’s point of contact with objects. We constructed the BeadSight sensor, which does not rely on any calibration and instead relies entirely on neural networks to distill information about contacts and movements on the tactile surface.
In this work, we repeated the visuo-tactile pre-training USB plugging experiment using the low-fidelity BeadSight sensor to directly compare the effectiveness of the different tactile sensors for training imitation learning agents on dexterous manipulation tasks. We also validate this approach with a second drawer pick-and-place experiment and investigate how using different datasets for visuo-tactile pre-training affects the agent’s performance on downstream tasks. Additionally, we confirm that tactile information is responsible for the improvement in agent performance by comparing with a vision-only pre-training approach.
Our contributions are:
@@
    We demonstrate the effectiveness of visuo-tactile pre-training with a low-fidelity, open-source sensor.
@@
    We demonstrate an affordable, durable approach to visuo-tactile pre-training, which is especially cost-effective for improving the performance of vision-only agents.
@@
    We show that by freezing a pre-trained tactile encoder or removing it entirely, we can mitigate the overfitting of low-fidelity tactile sensors while still maintaining a performance boost from tactile information.
  4. 4.
    We demonstrate the unique impact of tactile sensing on performance by comparing with a vision-only pre-training ablation.
II BACKGROUND AND RELATED WORK
II-A Tactile Sensing
In robotic manipulation, there are a wide variety of tactile sensors and sensing approaches [1]. These sensors use different modalities to collect tactile information, including forces and torques applied normally or tangentially [5] [6], mechanical vibrations [7] [8], thermal measurement and conductivity [9] [10], and pretouch proximity [11].
@@ II-A Tactile Sensing
BeadSight is a low-cost open-source visuo-tactile sensor that uses hydrogel beads as an inexpensive, durable sensing medium. An embedded camera observes the beads’ motion and deformation, however, the beads’ movements are stochastic and, unlike GelSight, the relationship between the beads and shear forces are not directly computed from first-principle physics methods. Previously, a deep neural network was used to reconstruct contact pressure maps with BeadSight as an example tactile learning approach [4].
II-B Tactile Sensing and Manipulation Control Policies
Prior work integrating tactile sensors in control policies range from classical control approaches to reinforcement and imitation learning. [16] used classical state estimation and a heuristic method to follow cables and insert wires. [17] combined a linear-quadratic regulator (LQR) controller and linear dynamics model to follow an audio cable and successfully plug in a headphone jack. Reinforcement Learning (RL) agents have been trained to perform peg insertion, door opening, and in-hand rotation tasks [18, 19]. However, these approaches require simulation for the sheer amount of training required ($\sim$ 1 million steps). Nearest Neighbors Imitation Learning [20] is a method that encodes observations and demonstrations into a latent space and compares the distance between them; [21] used this method with online residual reinforcement learning to reduce computation intensity and learn visuo-tactile peg insertion. [22] used a similar approach to learn a variety of tasks including bowl and cup unstacking, bottle opening, and joystick movement.
II-C Imitation Learning Methods
Imitation Learning (IL) methods aim to mimic human behavior by learning from human demonstrations, which can be easier than designing optimal reward functions as in reinforcement learning [23]. Variations of this method have trained manipulation agents to perform kitchen tasks and stack blocks [24, 25] and even sculpt clay [26]. IL tends to perform more poorly for more complex action sequences, particularly with non-deterministic goal policies and significant shifts between training and deployment domains. A state-of-the-art method was selected that aims to address these drawbacks, Diffusion Policy [27], which uses diffusion to better model multi-modal action pathways.
II-C1 Diffusion Policy
@@
II-C2 Action Chunking Transformers (ACT)
@@
ACT trains a Conditional Variational Autoencoder (CVAE) combined with a transformer backbone to predict action sequences conditioned on state and vision observations [30]. During training, the latent variable of the CVAE helps the model capture multi-modal policies, while a large KL divergence loss keeps the network from becoming too reliant on the latent information. During inference, the latent variable is set to zero. The autoencoder uses a latent variable to reduce the impact of multi-modal training data, and temporal ensembling at inference reduces the effect of a single bad prediction. Specifically, ACT predicts the goal states $A_{t,t}$ to $A_{t,t+h}$ at each timestep. During inference, all predictions $A_{t-h,t}$ to $A_{t,t}$ of the goal action at the current timestep are combined using a weighted average of exponentially decaying weights, $w_{i}=e^{-ki}$, and the ensembled action is executed.
II-D Contrastive Pre-training
@@
Contrastive pre-training methods train models to distinguish between similar and dissimilar data, coding them as positive (similar) or negative (dissimilar) pairs [31]. Those modeled after the structure of Contrastive Language Image Pre-training (CLIP) [32] do this by maximizing the cosine similarity of the paired data in some multimodal embedding space, while minimizing the similarity of dissimilar pairs. For visuo-tactile robotic manipulation, tactile and visual observations can be compared in that latent space. [33] used this approach to better identify flaws in fabric, and [34, 35] both improved visuo-tactile object identification this way.
III METHODS
Figure 2: BeadSight components and construction, including (from left to right) the camera, housing, lid frame, acrylic panel, and bead sac on a light-blocking backing
Figure 3: The embedded camera’s view of the BeadSight hydro-gel bead sac, relaxed (1) and pressed (2), after down-scaling and correcting fish eye distortion.
III-A BeadSight Hardware
@@
The BeadSight sensor used in this work measures contact forces using the deformation of hydro-gel beads, captured by a rear-mounted camera. Its 3D-printed enclosure (Figure 2) includes threaded heat inserts for attachments and a mount for the Franka Emika Panda robot gripper. The housing and lid allow easy replacement of the contact surface, enhancing durability. The lid consists of a 3D-printed frame, a press-fit acrylic insert, and a hydro-gel bead sac, which is backed with white adhesive tape to improve light reflection. Black PVC adhesive tape seals the sensor from external light sources (Figure 3).
@@
Using the same method as [4], the hydro-gel bead sacs use Polyacrylamide (PAM) beads hydrated with water and sealed in clear PVC sheeting. For this work, each sac contained 50 green PAM beads hydrated with 5ml of water. The embedded camera, also from [4], is a 1080 × 1920 USB camera with a 180° fish-eye lens, capturing at 30Hz. Four LEDs illuminate the bead sac, and images are downscaled to 480 × 480 resolution with fish-eye correction before processing.
III-B Data Collection
We selected two primary tasks to evaluate the low-fidelity sensor and pre-training method, using the same observation sensors in both cases. For each task, 100 expert demonstrations were collected via teleoperation of a Franka Emika Panda robot. Teleoperation was performed using the Oculus VR method from [36], where handheld Oculus controllers controlled the robot’s end-effector, while the operator viewed the workspace directly rather than through the VR headset.
@@ III-B Data Collection
Visual observations were captured by six Intel RealSense cameras: four D415 cameras around the workspace, a D415 wrist-mounted camera, and a D445 overview camera (Figure 4). The BeadSight sensor on the robot’s right finger collected tactile data, with a 3D-printed left finger featuring a silicone gel surface to aid grip.
Figure 4: USB cable plugging experimental scene, with a D415 wrist-mounted camera (not visible) positioned behind the robot gripper.
Figure 5: Drawer placement task experimental scene, including a side view of the BeadSight Sensor attached to the end effector.
The first task, USB cable plugging (Figure 4), uses the same experimental setup as [3] to allow for comparison with the GelSight sensor. This dexterous, precision-based task relies heavily on tactile information. For Diffusion Policy, we tested both ablating and freezing the tactile encoder during the downstream task.
The second task is a longer-horizon sequence: opening a drawer, picking up a block, placing it inside, and then closing the drawer (Figure 5). Additionally, we introduced a supportive pre-training task (picking and placing the same block onto the drawer handle) to assess the impact of pre-training with identical, similar, and dissimilar tasks on downstream performance.
III-C Tactile Encoder
@@
Because the BeadSight sensor’s hydro-gel beads are not fixed, tactile events are inferred from the beads’ relative motion and deformation over time. To capture this relative motion, we introduce a “tactile horizon”, the total number of past tactile frames (including the current frame) combined to form the tactile observation. We use a tactile horizon of 5 frames (the current time step and 4 previous time steps). We collapse the 5 frames on the channel dimension, converting 5 individual $3\times 480\times 480$ images to a single $15\times 480\times 480$ tensor. For a different tactile horizon, the representative tensor would be of the shape $3h\times 480\times 480$, where $h$ is the horizon. The tactile observation is processed by a tactile encoder consisting of a pre-trained ResNet-18 [37] CNN (with the final classification layers removed), combined with a single convolutional layer (to map the $3h$ channels to 3 channels), resulting in a $\mathbb{R}^{3\times h\times 480\times 480}\to\mathbb{R}^{512}$ embedding.
@@
To provide a horizon for the first few time steps, the BeadSight’s tactile observation buffer is initialized with $h-1$ duplicates of the initial tactile frame. For these tasks, the sensor is not in contact with any object or surface at the start of each task run, so we would expect the initial frames to be nearly identical.
III-C1 Contrastive pre-training
@@
The vision and tactile encoders were pre-trained using the contrastive pre-training method described in [3]. This method leverages the relationships between the three different modalities of data our agent observes (visual, tactile, and positional) to train the encoders to extract task-relevant features. During pre-training, two projection networks are learned, one for visual observations and one for tactile and positional observations. These projection networks consist of an encoder (ResNet-18 for the image encoder and the BeadSight tactile encoder) along with a projection head that casts the encoded observations to a shared latent embedding space. The positional information is also passed to the tactile projection head. By adding position information to the tactile projector, we ensure the resulting latent space has both global information (position observations) and local information (tactile observations).
@@
To train the projection networks, a set of timesteps is sampled from a single trajectory, and the corresponding observations are projected to the shared latent space. Then, a CLIP loss is used to encourage the embeddings from the same timestep to be similar, while driving apart the projection of the observations from different timesteps. All of the camera views are encoded using a shared visual encoder. At each update, the contrastive loss is calculated separately for each camera and then summed, according to the equation:
$loss=-\sum_{c}\sum_{i\in t}\frac{1}{2n}(\log(\frac{\exp(sim_{i,i,c}/\tau)}{\sum_{j\in t}\exp(sim_{i,j,c}/\tau)})$
$+\log(\frac{\exp(sim_{i,i,c}/\tau)}{\sum_{j\in t}\exp(sim_{j,i,c}/\tau)}))$
@@
III-C2 Diffusion Policy
Our Diffusion Policy approach leveraged [27], and was modeled after [3], with similar baseline hyper-parameters. We used the 1-D temporal CNN implementation, which was noted to be less sensitive to hyper-parameter tuning and to train stably for most tasks. Diffusion Policy is most effective when each camera view uses a separate trained encoder, so 6 ResNet-18 encoders were constructed and fine-tuned during training, one for each scene camera. For the pretrained model, six copies of the pretrained ResNet-18 vision encoder were constructed and separately fine-tuned during training.
After encoding, the vision and tactile observations were stacked to form a single observation vector, then passed to the network to condition its noise prediction. Our implementation uses an observation horizon of 1, and an action prediction horizon of 20. We also took advantage of noise scheduler decoupling, using 100 denoising steps during training for better resolution but using only 10 inference steps for faster execution.
III-C3 Action Chunking Transformers